Find dy dx if y = In(x2/x + 5). dy dxFind dp da if p = In 91 9 in(92-5). dp bpFind if p = In(94 + + 5). da dp da II
Answer #1

de [In (x? Vx+5)] 1 = d dx VX + 5 x? V x + 5 de [22] · Væ+5+ 22. d dx [Vx+5 22 x + 5 -1 2xVx+ 5 + 22- ] (x+5)7-1. [2x + 5] x2q2 – 5 d da [L( * a ज 9 q? = d dq q2 – 5 q de [e?-5].9-(02-5). dia a a q2 q2 – 5 (, [92] + (-5)) 9-1(92 - 5) q (q2 – 5) = -q2dla (ln(94 +5)] dq 1 = q4 + 5 . [24+5 d da 94] + ., [5] = 94 + 5 4q3 + 0 = 94 +5 4q3 94 + 5